OpenStudy (kaioken360):

f(x)=4x^3 and g(x)=2x. What is f(x)/g(x)? A) 4x B) 2x^2 C) 2x^3 D) 8x^4

OpenStudy (anonymous):

put \(f\) over \(g\) and then cancel some stuff

OpenStudy (anonymous):

\[f(x)=4x^3~and~g(x)=2x\]substitute it into\[\frac{ f(x) }{ g(x) }\]\[\frac{ f(x) }{ g(x) }=\frac{ 4x^3 }{ 2x }\]simplify it

OpenStudy (anonymous):

we noticed that 4x^3 and 2x can be divide by 2x

OpenStudy (anonymous):

\[=\frac{ 2x^2\times2x }{ 1\times2x }\]

OpenStudy (anonymous):

\[=\frac{ 2x^2\times\cancel{2x} }{ 1\times\cancel{2x} }\]

OpenStudy (anonymous):

\[=?\]solve it @kaioken360